OpenStudy (anonymous):

Write the equation of the line that passes through (–2, 6) and (2, 14) in slope-intercept form

OpenStudy (anonymous):

First, put your points in a coordinate axis  operating a=8 b=4 the form of your equation is: \[y=mx+i\] where m:slope i:Intercept so you need the slope and the intercept you can find the slope with this expression \[m=tg \alpha=\frac{ a }{ b }=\frac{ 8 }{ 4 }=2\] this is the equation of your line \[y=2x+i\] and you have two points of this. Replace one of the points in your equation for (-2,6): x=-2 y=6 \[6=2(-2)+i\] \[10=i\] \[y=2x+10\]

OpenStudy (anonymous):

y = 2x + 10 correct answer

OpenStudy (anonymous):

y-6 = {(14-6)/[2-(-2)]} x {x-(-2)} y-6 = 2(x+2) y = 2x +10 <<y=mx+c form>>